A data set lists weights​ (lb) of plastic discarded by households. The highest weight is 5.15 ​lb, the mean of all of the weights is x overbarequals1.888 ​lb, and the standard deviation of the weights is sequals2.073 lb. a. What is the difference between the weight of 5.15 lb and the mean of the​ weights? b. How many standard deviations is that​ [the difference found in part​ (a)]? c. Convert the weight of 5.15 lb to a z score. d. If we consider weights that convert to z scores between minus2 and 2 to be neither significantly low nor significantly​ high, is the weight of 5.15 lb​ significant?
Answer #1

Mean = 1.888 lb

Standard deviation = 2.073

a) Difference = 5.15 - 1.888 = 3.262

b) How many standard deviation = 3.262/2.073 = 1.57 (approx), rounded to 2 decimal places)

c) z-score = 1.57 (because z-score is equal to how many standard deviation far from mean)

d) Since z-score for 5.15 lb is between -2 and 2 hence weight 5.15 lb is not significant.

Please comment if any doubt. Thank you.